OpenStudy (anonymous):

Use the function defined below to find an equation of the tangent line at the point indicated. y=sinx, x=pie/3

OpenStudy (anonymous):

i have my slope of the tangent = root3/2

OpenStudy (anonymous):

hey again y' = cos(x) y'(pie/3) = cos(pie/3) = 1/2

OpenStudy (anonymous):

so thats my slope

OpenStudy (anonymous):

yes.. 1/2 is the slope

OpenStudy (anonymous):

now use y=y1+m(x-x1) in order to find the tangent line

OpenStudy (anonymous):

is my point (pie/3, sin(1/2)) ??

OpenStudy (anonymous):

no.. (pie/3,sin(pie/3))

OpenStudy (anonymous):

and we should write pi :P

OpenStudy (anonymous):

i need to write it in the form y=

OpenStudy (anonymous):

use this .. now you have a point (x1,y1) and a slope m y=y1+m(x-x1)

OpenStudy (anonymous):

\[y-\sin \frac{ \pi }{ 3 }=\frac{ 1 }{ 2 }x-\frac{ \pi }{ 6 }\]

OpenStudy (anonymous):

good..

OpenStudy (anonymous):

Sin(pi/3) = sqrt(3)/2

OpenStudy (anonymous):

so you can arrange it as : y(x) = x/2 +sqrt(3)/2 - Pi/6 or y(x) = x/2 + (3sqrt(3)-Pi)/6

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw